# Stochastic heat flow by moments

### v1

@@ 1.1. Corollaries and future directions
Let $Z^{\theta}$ be the SHF($\theta$). Then for every $u\in\mathbb{R}$, $z\in\mathbb{R}^{2}$, and $r>0$, $Z^{\theta}_{\cdot\,+u,\,\cdot\,+u}=Z^{\theta}$ in law, $Z^{\theta}\circ\tau_{z}^{\otimes 2}=Z^{\theta}$ in law, and $Z^{\theta}_{r\,\cdot,r\,\cdot}\circ\sigma^{2}_{\sqrt{r}}{}^{\otimes 2}=rZ^{\theta+\log r}$ in law.
@@
$\sup\Big{\{}\Big{|}\mathbf{E}\Big{[}\prod_{i=1}^{n}\big{(}Z^{\varepsilon}_{s,t}\,g^{\prime}_{i}\otimes g_{i}\big{)}\Big{]}-\Big{\langle}\bigotimes_{i=1}^{n}g^{\prime}_{i},\mathcal{Q}^{[\![n]\!]}(t-s)\bigotimes_{i=1}^{n}g_{i}\Big{\rangle}\Big{|}\ \Big{|}\ \|g^{\prime}_{i}\|_{L^{2}(\mathbb{R}^{2})}\leq 1\Big{\}}\xrightarrow{\varepsilon\to 0}0\ .$ (1.7)
Together, Propositions 1.4 and 1.6 immediately give the corollary.
@@
The SHF($\theta$) satisfies Definition 1.2(d) for all $\overline{n}<\infty$.
A potential future direction is to use Theorem 1.3 to prove the conjecture of Jeremy Clark on the conditional Gaussian Multiplicative Chaos (GMC) property. First, the recent work [CM24] constructed the polymer measure associated with the SHF; see Section 1.3 for more discussion. Given the polymer measure, the conjecture states that, if one takes the polymer measure with parameter $\theta$ as the base measure (which is random) and construct from it the GMC of strength $\eta>0$, the result should be the polymer measure with parameter $\theta+\eta$. Based on [CM23], the work [CM24] showed that the second moment behaves consistently with this conjecture and outlined a path toward the conjecture (in Section 2.6 there) assuming “a list of properties uniquely characterizing the distribution of the 2d SHF”. Theorem 1.3 seems applicable for this purpose if one can push the consistency up to the fourth moment.
Another potential direction is to study the SHF as a noise. By Definition 1.2 and Corollary 1.5, the SHE should generate a noise in the sense of [TV98, Tsi04]. An interesting question asks whether the noise is black in the sense of [TV98, Tsi04]. It has be recently proven in [HP24] that the directed landscape, which the large-time limit of the logarithm of the one-dimensional SHE [Wu23] (see also [QS23, Vir20]), is a black noise. On the other hand, the one-dimensional SHE itself should be a white noise, again in the sense of [TV98, Tsi04], because the solution is measurable with respect to the driving spacetime white noise.
1.2. Example: Geometric Brownian Flow
Here we use the Geometric Brownian Flow (GBF, defined below) as an elementary example to illustrate how, in the presence of fast moment growth, one can still use the moments to uniquely characterize the law of the flow. The discussion here also serves as a good overview of the paper, because the proof of the GBF largely parallels that of the SHF. To highlight the similarities between the GBF and SHF, slightly abusing notation, we also use $Z,\widetilde{Z}$ to denote the GBF.
Definition 1.8.
We call $Z$ a GBF, if
  1. (a)
    $Z=Z_{s,t}$ is a $(0,\infty)$-valued, continuous process on $\mathbb{R}^{2}_{\scriptscriptstyle\leq}:=\{(s\leq t)\in\mathbb{R}^{2}\}$,
  2. (b)
    for all $s<t<u$, $Z_{s,t}Z_{t,u}=Z_{s,u}$,
  3. (c)
    for all $s<t<u$, $Z_{s,t}$ and $Z_{t,u}$ are independent, and
  4. (d)
    for $n=1,\ldots,4$ and $s<t$, $\mathbf{E}[Z_{s,t}^{n}]=e^{(t-s)n^{2}/2}$.
Indeed, $Z_{s,t}=e^{B(t)-B(s)}$, where $B$ is the two-sided Brownian motion, satisfies Definition 1.8. For fixed $s<t$, $e^{B(t)-B(s)}$ is log-normal and has an indeterminate moment problem. Still, the first four moments over all time intervals uniquely characterize the law of the flow.
Proposition 1.9.
If $Z$ and $\widetilde{Z}$ are GBF, then $Z$ and $\widetilde{Z}$ are equal in law.
Proof.
By Definition 1.8(a)(c), it suffices to prove that, for every $s<t$, $Z_{s,t}=\widetilde{Z}_{s,t}$ in law. Take $(s,t)=(0,1)$ to simplify notation and take any $h\in C^{4}(\mathbb{R})$ with bounded derivatives. Our goal is to prove that $\mathbf{E}[h(Z_{0,1})]-\mathbf{E}[h(\widetilde{Z}_{0,1})]=0$. Use Lindeberg’s exchange method and write
$\mathbf{E}[h(Z_{0,1})]-\mathbf{E}[h(\widetilde{Z}_{0,1})]=\sum_{i=1}^{N}\big{(}\mathbf{E}[h(\widetilde{Z}_{0,s}Z_{s,1})]-\mathbf{E}[h(\widetilde{Z}_{0,t}Z_{t,1})]\big{)}\big{|}_{s,t=(i-1)/N,i/N}\ .$ (1.8)
Hereafter, we take $Z$ and $\widetilde{Z}$ to be independent within expectations like the ones above. Next, write $W:=Z-1$ and $\widetilde{W}:=\widetilde{Z}-1$ and use Definition 1.8(d) to get
$\mathbf{E}[W^{n}_{s,t}]=\mathbf{E}[\widetilde{W}^{n}_{s,t}]=\sum_{k=0}^{n}\binom{n}{k}(-1)^{n-k}e^{k^{2}(t-s)/2}=:w^{n}_{t-s}\ ,\qquad n=1,\ldots,4\ .$ (1.9)
Consider the first expectation in (1.8). Use Definition 1.8(b) to write $Z_{s,1}=Z_{s,t}Z_{t,1}$ and write $Z_{s,t}=1+W_{s,t}$. This way $\widetilde{Z}_{0,s}Z_{s,1}=Y_{s,t}+V_{s,t}$, where $Y_{s,t}:=\widetilde{Z}_{0,s}Z_{t,1}$, and $V_{s,t}:=\widetilde{Z}_{0,s}W_{s,t}Z_{t,1}$. Taylor expand $h$ up to the third order to get
$\mathbf{E}[h(\widetilde{Z}_{0,s}Z_{s,1})]=\sum_{n=0}^{3}\frac{1}{n!}\mathbf{E}\big{[}h^{(n)}\big{(}Y_{s,t}\big{)}\cdot V_{s,t}^{n}\big{]}+v_{4}\ ,$ (1.10)
where $|v_{4}|\leq c(h)\,\mathbf{E}[V_{s,t}^{4}]$, and $c(h)$ is a constant that depends only on $h$. For the $n\leq 3$ terms, write the expectation as $\mathbf{E}[\mathbf{E}[\,\cdot\,|\widetilde{Z}_{0,s},Z_{t,1}]]$ and evaluate the inner expectation using the independence of $\widetilde{Z}_{0,s},W_{s,t},\widetilde{Z}_{t,1}$ and (1.9) to get
$\mathbf{E}\big{[}h^{(n)}\big{(}Y_{s,t}\big{)}\cdot V_{s,t}^{n}\big{]}=\mathbf{E}\big{[}h^{(n)}(\widetilde{Z}_{0,s}Z_{t,1})\cdot\widetilde{Z}_{0,s}^{n}\cdot w^{n}_{s,t}\cdot Z_{t,1}^{n}\big{]}=:h^{n}_{s,t}\ ,\qquad n\leq 3\ .$ (1.11)
Similarly, $|v_{4}|\leq c(h)\mathbf{E}[Z_{0,s}^{4}]w^{4}_{s,t}\mathbf{E}[Z_{t,1}^{4}]\leq c(h)w^{4}_{t-s}$. It is straightforward to check from (1.9) that $w^{4}_{t-s}\leq c\,(t-s)^{2}$, for all $s<t\in[0,1]$ and some universal constant $c$. Hence
$\Big{|}\mathbf{E}[h(\widetilde{Z}_{0,s}Z_{s,1})]-\sum_{n=0}^{3}\frac{1}{n!}\mathbf{E}\big{[}h^{(n)}(\widetilde{Z}_{0,s}Z_{t,1})\cdot\widetilde{Z}_{0,s}^{n}\cdot w^{n}_{s,t}\cdot Z_{t,1}^{n}\big{]}\Big{|}\leq c(h)\,(t-s)^{2}\ .$ (1.12)
Next, apply the same argument above to $\mathbf{E}[h(\widetilde{Z}_{0,t}Z_{t,1})]=\mathbf{E}[h(\widetilde{Z}_{0,s}(1+\widetilde{W}_{s,t})Z_{t,1})]$. The result shows that (1.12) holds for $\mathbf{E}[h(\widetilde{Z}_{0,t}Z_{t,1})]$ in place of $\mathbf{E}[h(\widetilde{Z}_{0,s}Z_{s,1})]$. Combining (1.12) for $\mathbf{E}[h(\widetilde{Z}_{0,s}Z_{s,1})]$ and for $\mathbf{E}[h(\widetilde{Z}_{0,t}Z_{t,1})]$ and inserting the result into (1.8) give
$\big{|}\mathbf{E}[h(Z_{0,1})]-\mathbf{E}[h(\widetilde{Z}_{0,1})]\big{|}\leq c(h)\,N\cdot\big{(}\tfrac{1}{N}\big{)}^{2}\ .$ (1.13)
Sending $N\to\infty$ completes the proof. ∎
Our proof of Theorem 1.3 largely parallels that of Proposition 1.9. One difference is that the former requires working with the product in Definition 1.1, while the latter requires only the usual product of real numbers. Another difference is that obtaining the analog of the bound $|v^{4}_{t-s}|\leq c(h)(t-s)^{2}$ requires much more work, which will be carried out in Section 3.2.
1.3. literature
@@ 1.2. Example: Geometric Brownian Flow
An important tool in studying the SHF is the two-dimensional delta-Bose gas, which is a quantum many-body system with an attractive, pairwise delta potential. Quantum systems with delta potentials exhibit intriguing physical properties, post significant mathematical challenges, and enjoy a long history of study; see [AGHKH88] for a review. We mention the works [DFT94, Raj99, DR04, GHL20, GH22] on the two-dimensional delta-Bose gas. On the SHF side, based on [AGHKH88, ABD95], the work [BC98] established the convergence of the second moment of the noise-mollified SHE. The work [Fen15] carried out detailed analysis of the second moment and further showed that the pointwise limits of the third and second moments behave inconsistently with the lognormal distribution. Based on techniques developed in [CSZ19a] to control the chaos series, the work [CSZ19b] obtained the convergence of the third moment of the discrete polymers and mollified SHE. Later, the work [GQT21] obtained the convergence of all moments of the mollified SHE. This is done based on the works [Raj99, DR04], which give a transparent and elegant description of the resolvent of the delta-Bose gas. These works used a crucial $L^{2}$ bound from [DFT94, Lemma 3.1]. The $L^{p}$ analog of the bound was later proven and used in [CSZ23a, Lemma 6.8], and the current paper also uses this $L^{p}$ bound; see (2.9c) and the comment that follows. The works [Che22, Che24a, Che24b] develop probabilistic approaches to study the delta-Bose gas. High moments of the polymers and the mollified SHE are studied in [CZ23, LZ23, CZ24].
Closely related to the SHF is its polymer measure. In one dimension, the work [AKQ14] constructed the polymer measure from the SHE by using the Chapman–Kolmogorov equation. The idea of [AKQ14] could generalize to the SHF, but since the SHF is measure-valued rather than function-valued, formulating and proving the Chapman–Kolmogorov equation are subtle and significantly more challenging. The recent work [CM24] achieved these and constructed the polymer measure from a limit of the mollified SHE. In Section 2.2, we do similarly but from $Z$ (that satisfies Definition 1.2) and for the finite-dimensional distributions, as an intermediate step in our proof of Theorem 1.3.
An interesting question asks whether the polymer measure is a GMC. For the two-dimensional SHE/parabolic Anderson model driven by the spatial white noise, the work [QRV22] showed that the polymer measure is a GMC. The same method should apply to the one-dimensional SHE (driven by the spacetime white noise). For the SHF, the work [CSZ23b] showed that the endpoint distributions of the polymer measure is not a GMC, and the work [CM24] explained in Section 2.6 that the entire polymer measure is not a GMC with the Brownian motion being its based measure. On the other hand, Clark conjectured the conditional GMC property mentioned in Section 1.1. The analogous property was proven for the polymer measure on the diamond hierarchical lattice [Cla23], which is a model that exhibits similar behaviors as the SHF and has been studied in [ACK17, AC19, Cla19, Cla22, CL23].
@@
At sub- and quasi-critical temperatures, Gaussian limits are obtained for the polymers and the mollified SHE [CSZ17, CD20, CSZ20, Gu20, CCR23, NN23] and for the nonlinear mollified SHEs [Tao24, DG24]. The limiting behaviors of the one-point distributions have been studied in [CSZ17, DG22, DG23]. In dimension three and higher, the law of large numbers and fluctuations of the polymers, mollified SHE, and its nonlinear counterparts have been studied in [MU18, GRZ18, DGRZ20, CCM20, CCM24, MSZ16, CL17, CCM20, CN21, CNN22, CCM24, DM24].
We mentioned some recent works on the two-dimensional anisotropic KPZ (aKPZ) equation. Under a weak scaling that logarithmically attenuates the nonlinearity, [CET23a, CES21] obtained the tightness and Gaussian limit of the aKPZ equation. Later, [CET20, CET23b] showed that, without the weak scaling, the aKPZ exhibits a surprising logarithmic superdiffusive behavior.
@@ Acknowledgment
I thank Yu Gu, Jeremy Quastel, and Bálint Virág for many useful and inspiring discussions. This research is partially supported by the NSF through DMS-2243112 and the Alfred P. Sloan Foundation through the Sloan Research Fellowship FG-2022-19308.
@@
In Section 2, we prepare the notation and tools. In Section 3, we prove Theorem 1.3. In Section 4, we prove Proposition 1.4. The proof of some technical results are placed in the appendix.
@@ 2.1. Delta-Bose semigroup
$\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha}$ $:=\mathbb{R}^{2}\times\mathbb{R}^{2(\omega\setminus\alpha)}:=\big{\{}y=(y_{\mathrm{c}},(y_{i})_{i\in\omega\setminus\alpha})\,\big{|}\,y_{c},y_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (2.2)
@@
Let $\mathsf{p}(t,x):=\exp(-|x|^{2}/2t)/(2\pi t)$ denote the heat kernel on $\mathbb{R}^{2}$, where $x\in\mathbb{R}^{2}$ and ||$\cdot$|| ⋅ | denotes the Euclidean norm on $\mathbb{R}^{d}$. Let $\mathcal{P}^{\omega}(t,x):=\prod_{i\in\omega}\mathsf{p}(t,x_{i})$ denote the heat kernel on $\mathbb{R}^{2\omega}$, and let
@@
For $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}(\omega)$, define the integral operators $\mathcal{P}^{\omega}_{\alpha}(t)$, $\mathcal{P}^{\omega}_{\alpha}(t)^{*}$, $\mathcal{P}^{\omega}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}^{\omega}_{\alpha}(t)=\mathcal{J}^{\theta,\omega}_{\alpha}(t)$ through their kernels as
@@
$\mathcal{P}^{\omega}_{\alpha\alpha^{\prime}}(t,y,y^{\prime})$ $:=\mathcal{P}^{\omega}\big{(}t,S_{\alpha}y-S_{\alpha^{\prime}}y^{\prime}\big{)}\ ,$ $y\in\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha}\ ,y^{\prime}\in\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha^{\prime}}\ ,$ (2.5b)
@@
where $11/p^{\prime}+1/q^{\prime}=$ 1. We will mostly omit the underlying spaces $\Omega$ when writing the norms, because the spaces can be read off from the definition of the operators. For example, referring to (2.5a), we see that $\|\mathcal{P}^{\omega}_{\alpha}(t)\|_{p,a\to p^{\prime},a^{\prime}}=\|\mathcal{P}^{\omega}_{\alpha}(t)\|_{L^{p,a}(\mathbb{R}^{2\omega})\to L^{p^{\prime},a^{\prime}}(\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha})}$. Next, note that the kernels in (2.5a)–(2.5c) are nonnegative. For a one-parameter family of integral operators $\mathcal{T}(t)$, $t>0$, with a nonnegative kernel $\mathcal{T}(t,z,z^{\prime})$, similarly define
@@
where the supremum runs over $\|f^{\prime}\|_{q^{\prime},a}\leq 1$ and $\|f\|_{p,a}\leq 1$, with $11/p^{\prime}+1/q^{\prime}=$ 1. If this quantity is finite, the operator $\int_{0}^{b}\mathrm{d}t\,\mathcal{T}(t)$ is well-defined and bounded. When $a=0$, the weighted norm $\|\cdot\|_{p,a}$ reduces to the usual $L^{p}$ norm, which we write as $\|\cdot\|_{p,0}=\|\cdot\|_{p}$. Hereafter, write $c=c(v_{1},v_{2},\ldots)$ for a general, finite, positive, deterministic constant that may change from place to place but depends only on the designated variables $v_{1},v_{2},\ldots$. Recall that $\theta$ is fixed throughout the paper, and we (often) drop the dependence on $\theta$. Hence every constant in this paper may implicitly depend on $\theta$. Given any $p\in(1,\infty)$, $a\in[0,\infty)$, and finite $\omega\subset\mathbb{Z}_{>0}$, there exists $c=c(|\omega|,p,a)=c(\theta,|\omega|,p,a)$ such that for all $t>0$ and $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}(\omega)$ and $11/p+1/q=$ 1,
@@
We give a proof of (2.9) in the appendix based on [GQT21], [CSZ23a, Lemma 6.8], and [ST24]. In particular, (2.9c) essentially follows from [CSZ23a, Lemma 6.8], and the comparison argument from [ST24] reduces (2.9) to its $a=0$ counterpart. Let us state a few more technical bounds, which will be used later and are also proven in the appendix. With $11/p+1/q=$ 1 and $\alpha\in\mathrm{Pair}(\omega)$,
@@
This set indexes certain diagrams, hence the name $\mathrm{Dgm}$; see [GQT21, Section 2]. Write $|\vec{\alpha}|:=m$ for the length of $\vec{\alpha}\in\mathrm{Dgm}(\omega)$ and let $[\![n]\!]:=\{1,\ldots,n\}$. For $f=f(\tau,\tau^{\prime},\tau^{\prime\prime},\ldots)$ that depends on finitely many nonnegative $\tau$s, write $\int_{\Sigma(t)}\mathrm{d}\vec{\tau}f=\int_{\tau+\tau^{\prime}+\ldots=t}\mathrm{d}\vec{\tau}f$ for the convolution-like integral. For $\vec{\alpha}\in\mathrm{Dgm}(\omega)$, let
@@
Hereafter, products of operators are understood in the written order, so $\prod_{k=1}^{K}\mathcal{T}_{k}:=\mathcal{T}_{1}\mathcal{T}_{2}\cdots\mathcal{T}_{K}$. The delta-Bose semigroup on $\mathbb{R}^{2\omega}$ is
@@
To ensure the convergence of this sum, let us prepare a lemma. This lemma follows from the proof of [GQT21, Lemma 8.10], and we give a proof of it in the appendix for the sake of completeness.
@@
    $\Big{\|}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{2m+1}\mathcal{T}^{\prime}_{k/2}(\tau_{k/2})\Big{\|}_{\mathrm{op}}\leq\frac{c^{m}\,m^{3}\,t^{1+b+b^{\prime}}\,e^{\lambda t}}{(b+1)\wedge(b^{\prime}+1)\wedge 1}\Big{(}c_{1}^{m}+c_{0}^{2}\,\Big{(}c_{1}+\frac{c_{0}^{2}}{\log(\lambda-c_{0}-1)}\Big{)}^{m-1}\Big{)}\ ,$
@@
    $\Big{\|}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=2}^{2m+1}\mathcal{T}^{\prime}_{k/2}(\tau_{k/2})\Big{\|}_{\mathrm{op}}\leq\frac{c^{m}\,m^{3}\,t^{b^{\prime}}\,e^{\lambda t}}{(b^{\prime}+1)\wedge 1}\Big{(}c_{1}^{m}+c_{0}\,\Big{(}c_{1}+\frac{c_{0}^{2}}{\log(\lambda-c_{0}-1)}\Big{)}^{m-1}\Big{)}\ .$

### v3

@@ 1.4. Corollaries, generalizations
Let $Z^{\theta}$ be SHF($\theta$). Then for every $u\in\mathbb{R}$, $z\in\mathbb{R}^{2}$, and $r>0$, $Z^{\theta}_{\cdot\,+u,\,\cdot\,+u}=Z^{\theta}$ in law, $Z^{\theta}\circ\tau_{z}^{\otimes 2}=Z^{\theta}$ in law, and $Z^{\theta}_{r\,\cdot,r\,\cdot}\circ\sigma^{2}_{\sqrt{r}}{}^{\otimes 2}=rZ^{\theta+\log r}$ in law.
@@
$\sup\Big\{\Big|\mathbf{E}\Big[\prod_{i=1}^{n}Z^{\theta,\varepsilon}_{s,t}\,g^{\prime}_{i}\otimes g_{i}\Big]-\Big\langle\bigotimes_{i=1}^{n}g^{\prime}_{i},\mathcal{Q}^{[\![n]\!]}(t-s)\bigotimes_{i=1}^{n}g_{i}\Big\rangle\Big|\ \Big|\,\|g^{\prime}_{i}\|_{L^{2}(\mathbb{R}^{2})}\leq 1\Big\}\xrightarrow{\varepsilon\to 0}0\ .$ (1.10)
Together, Propositions 1.5 and 1.7 immediately give the corollary.
@@
SHF($\theta$) satisfies Definition 1.2(d) for all $\overline{n}<\infty$.
Next, we note that the choice of the test functions in Definition 1.2(d) has much flexibility. As we will see in Section 2.1, the delta-Bose semigroup $\mathcal{Q}^{[\![n]\!],\theta}(t)$ has as a positive kernel $\mathcal{Q}^{[\![n]\!],\theta}(t;x^{\prime},x)$, where $x,x^{\prime}\in\mathbb{R}^{2[\![n]\!]}$ and we write $n]\!]=\{1,\ldots,n\}$ and $\mathbb{R}^{2n}=\mathbb{R}^{2[\![n]\!]}$ to streamline subsequent notation. Hence, (1.2) is equivalent to
$\mathbf{E}Z_{s,t}^{\otimes n}=\mathrm{d}x\otimes\mathrm{d}x^{\prime}\,\mathcal{Q}^{[\![n]\!],\theta}(t-s;x^{\prime},x)\quad\text{in }\mathcal{M}_{+}(\mathbb{R}^{2[\![n]\!]}\times\mathbb{R}^{2[\![n]\!]})\ .$ (1.11)
We choose $L^{2}(\mathbb{R}^{2})^{\otimes n}$ to be the set of test functions just for convenience, and the choice can be any that uniquely characterizes measures in $\mathcal{M}_{+}(\mathbb{R}^{2[\![n]\!]}\times\mathbb{R}^{2[\![n]\!]})$.
We end this subsection by stating a slightly more general version of Theorem 1.3, which is the one that we will actually prove.
Theorem 1.9.
Take any deterministic dense $G\subset\mathbb{R}$ and consider $\mathcal{M}_{+}(\mathbb{R}^{4})$-valued processes $Z=Z_{s,t}$ and $\widetilde{Z}=\widetilde{Z}_{s,t}$ over $s<t\in G$. If $Z$ and $\widetilde{Z}$ satisfy Definition 1.2(b)(d) for $t_{0}<t_{1}<\ldots<t_{k}\in G$, then $Z$ and $\widetilde{Z}$ have the same finite dimensional distributions.
Indeed, by the continuity requirement in Definition 1.2(a), Theorem 1.9 for $G=\mathbb{R}$ immediately implies Theorem 1.3.
1.5. Some related literature
@@
An important tool in studying the SHF is the two-dimensional delta-Bose gas, which is a quantum many-body system with an attractive, pairwise delta potential. Quantum systems with delta potentials exhibit intriguing physical properties, pose significant mathematical challenges, and enjoy a long history of study; see [AGHKH88] for a review. We mention the works [DFT94, Raj99, DR04, GHL20, GH22] on the two-dimensional delta-Bose gas. On the SHF side, based on [AGHKH88, ABD95], the work [BC98] established the convergence of the second moment of the noise-mollified SHE. The work [Fen15] carried out detailed analysis of the second moment and further showed that the pointwise limits of the third and second moments behave inconsistently with the lognormal distribution. Based on techniques developed in [CSZ19a] to control the chaos series, the work [CSZ19b] obtained the convergence of the third moment of the discrete polymers and mollified SHE. Later, the work [GQT21] obtained the convergence of all moments of the mollified SHE. This is done based on the works [Raj99, DR04], which give a transparent and elegant description of the resolvent of the delta-Bose gas. These works used a crucial $L^{2}$ bound from [DFT94, Lemma 3.1]. The $L^{p}$ analog of the bound was later proven and used in [CSZ23a, Lemma 6.8], and the present paper also uses this $L^{p}$ bound; see (2.9c) and the comment that follows. The works [Che22, Che24a, Che24b] develop probabilistic approaches to study the delta-Bose gas. High moments of polymers are studied in [CZ23, CZ24, CN25], and real-power moments of the polymers in the subcritical regime are studied in [LZ23]. After the posting of the first version of this manuscript, [Par25b, Par25a] use the ideas of moments to show the convergence to the 1dd SHE of certain ill-behaved or non-Markov models.
@@
At sub- and quasi-critical temperatures, Gaussian limits are obtained for the polymers, the mollified SHE, and the mollified KPZ equation [CSZ17, CD20, CSZ20, Gu20, CCR23, NN23] and for the nonlinear mollified SHEs [Tao24, DG24]. Instances of the SHE and polymer with Lèvy noise have been studied [BL22, BCL23]. The limiting behaviors of the one-point distributions have been studied in [CSZ17, DG22, DG23, CD24]. In dimension three and higher, the law of large numbers and fluctuations of the polymers, mollified SHE, and its nonlinear counterparts have been studied in [MU18, GRZ18, DGRZ20, CCM20, CCM24, MSZ16, CL17, CCM20, CN21, CNN22, CCM24, DM24, Jun25]. There has been much study of the phase transition, scaling functions, and related behaviors of direct polymers with general weight. We refer to [Lac10, BL17, JL24, Lac25] and the references therein for some recent work.
@@ Acknowledgment
I thank Yu Gu, Jeremy Quastel, and Bálint Virág for many useful and inspiring discussions. I thank Jeremy Clark for many useful suggestions that improve the presentation of this paper, and I thank Maël Laoufi and Zuodi Xie for the careful reading of the manuscript. This research is partially supported by the NSF through DMS-2243112 and the Alfred P. Sloan Foundation through the Sloan Research Fellowship FG-2022-19308.
@@
In Section 2, we prepare the notation and tools. In Section 3, we prove Theorem 1.9, which immediately implies Theorem 1.3. In Section 4, we prove Proposition 1.5. The proof of some technical results are placed in the appendix.
@@ 2.1. Delta-Bose semigroup
$\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha}$ $:=\mathbb{R}^{2}\times\mathbb{R}^{2(\omega\setminus\alpha)}:=\big\{y=(y_{\mathrm{c}},(y_{i})_{i\in\omega\setminus\alpha})\,\big|\,y_{\mathrm{c}},y_{i}\in\mathbb{R}^{2}\big\}\ ,$ (2.2)
@@
Let $\mathsf{p}(t,x):=\exp(-|x|^{2}/2t)/(2\pi t)$ denote the heat kernel on $\mathbb{R}^{2}$, where $x\in\mathbb{R}^{2}$ and $|\cdot|$ denotes the Euclidean norm on $\mathbb{R}^{d}$. Let $\mathcal{P}^{\omega}(t,x):=\prod_{i\in\omega}\mathsf{p}(t,x_{i})$ denote the heat kernel on $\mathbb{R}^{2\omega}$, and let
@@
For $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}(\omega)$, define the integral operators $\mathcal{P}^{\omega}_{\alpha}(t)$, $\mathcal{P}^{\omega}_{\alpha}(t)^{*}$, $\mathcal{P}^{\omega}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}^{\omega}_{\alpha}(t)=\mathcal{J}^{\omega,\theta}_{\alpha}(t)$ through their kernels as
@@
$\mathcal{P}^{\omega}_{\alpha\alpha^{\prime}}(t,y,y^{\prime})$ $:=\mathcal{P}^{\omega}\big(t,S_{\alpha}y-S_{\alpha^{\prime}}y^{\prime}\big)\ ,$ $\hskip-40.0pty\in\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha}\ ,y^{\prime}\in\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha^{\prime}}\ ,$ (2.5b)
@@
where $1/p^{\prime}+1/q^{\prime}=1$. We will mostly omit the underlying spaces $\Omega$ when writing the norms, because the spaces can be read off from the definition of the operators. For example, referring to (2.5a), we see that $\|\mathcal{P}^{\omega}_{\alpha}(t)\|_{p,a\to p^{\prime},a^{\prime}}=\|\mathcal{P}^{\omega}_{\alpha}(t)\|_{L^{p,a}(\mathbb{R}^{2\omega})\to L^{p^{\prime},a^{\prime}}(\mathbb{R}^{2}\times\mathbb{R}^{2\omega\setminus\alpha})}$. Next, note that the kernels in (2.5a)–(2.5c) are positive. For a $t$-parameterized integral operator $\mathcal{T}(t)$ with a nonnegative kernel $\mathcal{T}(t,z^{\prime},z)$, define
@@
where the supremum runs over $\|f^{\prime}\|_{q^{\prime},a}\leq 1$ and $\|f\|_{p,a}\leq 1$, with $1/p^{\prime}+1/q^{\prime}=1$. If this quantity is finite, the operator $\int_{0}^{b}\mathrm{d}t\,\mathcal{T}(t)$ is well-defined and bounded. When $a=0$, the weighted norm $\|\cdot\|_{p,a}$ reduces to the usual $L^{p}$ norm, which we write as $\|\cdot\|_{p,0}=\|\cdot\|_{p}$. Hereafter, write $c=c(v_{1},v_{2},\ldots)$ for a general, finite, positive, deterministic constant that may change from place to place but depends only on the designated variables $v_{1},v_{2},\ldots$. Recall that $\theta$ is fixed throughout the paper, and we (often) drop the dependence on $\theta$. Hence, every constant in this paper may implicitly depend on $\theta$. Given any $p\in(1,\infty)$, $a\in[0,\infty)$, and finite $\omega\subset\mathbb{Z}_{>0}$, there exists $c=c(|\omega|,p,a)=c(\theta,|\omega|,p,a)$ such that for all $t>0$ and $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}(\omega)$ and $1/p+1/q=1$,
@@
We give a proof of (2.9) in the appendix based on [GQT21], [CSZ23a, Lemma 6.8], and [ST25]. In particular, (2.9c) essentially follows from [CSZ23a, Lemma 6.8], and the comparison argument from [ST25] reduces (2.9) to its $a=0$ counterpart. Let us state a few more technical bounds, which will be used later and are also proven in the appendix. With $1/p+1/q=1$ and $\alpha\in\mathrm{Pair}(\omega)$,
@@
[…]
@@
Hereafter, products of operators are understood in the written order, so $\prod_{\ell=1}^{k}\mathcal{T}_{\ell}:=\mathcal{T}_{1}\mathcal{T}_{2}\cdots\mathcal{T}_{k}$. The delta-Bose semigroup on $\mathbb{R}^{2\omega}$ is
@@
Specializing to $\omega=[\![n]\!]:=\{1,\ldots,n\}$ gives $\mathcal{Q}^{[\![n]\!]}(t)=\mathcal{Q}^{[\![n]\!],\theta}(t)$.
To ensure the convergence of (2.13), let us prepare a lemma. This lemma follows from the proof of [GQT21, Lemma 8.10], and we give a proof in the appendix for the sake of completeness.
@@
    $\Big\|\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{2m+1}\mathcal{T}^{\prime}_{k/2}(\tau_{k/2})\Big\|_{\mathrm{op}}\leq\frac{c^{m}\,c_{0}^{m+1}\,m^{3}\,t^{1+b+b^{\prime}}\,e^{\lambda t}}{(b+1)\wedge(b^{\prime}+1)\wedge 1}\Big(c_{1}^{m}+c_{0}\,\Big(c_{1}+\frac{c_{0}}{\log(\lambda-c_{0}-1)}\Big)^{m-1}\Big)\ ,$
@@
    $\Big\|\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=2}^{2m+1}\mathcal{T}^{\prime}_{k/2}(\tau_{k/2})\Big\|_{\mathrm{op}}\leq\frac{c^{m}\,c_{0}^{m}m^{3}\,t^{b^{\prime}}\,e^{\lambda t}}{(b^{\prime}+1)\wedge 1}\Big(c_{1}^{m}+c_{0}\,\Big(c_{1}+\frac{c_{0}}{\log(\lambda-c_{0}-1)}\Big)^{m-1}\Big)\ .$